Hi, please help. I tried doing this myself and ended up with x/x-1 as th e answer. Please show me the correct working.
1+1/x / 1-1/x^2
That is 1 plus 1 over x over(divided by) 1 minus 1 over x squared.

Answer by mukhopadhyay(490)

[1+1/x]/[1-(1/x)^2]
= [1+1/x]/[(1+1/x)(1-1/x)]
= 1/(1-1/x) = x/(x-1);